OpenStudy (anonymous):

i am having problems with integration of rational functions by partial fractions. the question is integral x+4/(x^2+5x-6)

OpenStudy (inkyvoyd):

so the first step of the partial fraction decomposition you are looking for is factorization of the denominator. Do you know how to do so?

OpenStudy (anonymous):

yes, what i have so far is \[X+4/(x+6)(x-1)\] which then you would have x+4=A(x-1)+B(x+6)

OpenStudy (inkyvoyd):

so as you can see you have a system of linear equations based on coefficient matching A*x+B*x=x -> A+B=1 A*(-1)+B*(6)=4 -> -A+6B=4 correct?

OpenStudy (anonymous):

ok i see what you did, but the answer i see in the back of the book is 1/7lnl(x+6)^2(x-1)^5l +C

OpenStudy (freckles):

do you want that method where you plug in values of x to solve for A and B?

OpenStudy (freckles):

\[\frac{x+4}{(x+6)(x-1)}=\frac{A}{x+6}+\frac{B}{x-1} \\ x+4=A(x-1)+B(x+6)\] well we know we can make one term 0 on the right by plugging 1 for x

OpenStudy (freckles):

so try to solve for B using x=1

OpenStudy (anonymous):

ok so B=5

OpenStudy (freckles):

let me check 1+4=A(1-1)+B(1+6) 5=A(0)+B(7) 5=0+7B 5=7B

OpenStudy (freckles):

B should be 5/7

OpenStudy (anonymous):

ok i didnt think you would use A for it

OpenStudy (freckles):

we didn't use A since (1-1)=0 and A(0)=0

OpenStudy (anonymous):

ok then, yeah i see

OpenStudy (inkyvoyd):

@freckles heaviside is awesome :o

OpenStudy (freckles):

ok so now you have \[\frac{x+4}{(x+6)(x-1)}=\frac{A}{x+6}+\frac{B}{x-1} \\ x+4=A(x-1)+B(x+6) \\ \text{ where } B=\frac{5}{7} \text{ So now we have } \\ x+4=A(x-1)+\frac{5}{7} (x+6)\] now to make that one term on the right 0 replace x with -6 instead Actually @inkyvovd I normally do the systems of equations way but I don't think op has learned that way

OpenStudy (anonymous):

so would A=1

OpenStudy (freckles):

let's find out... replacing x with -6... \[x+4=A(x-1)+\frac{5}{7}(x+6) \\ -6+4=A(-6-1)+\frac{5}{7}(-6+6) \\ -6+4=A(-6-1)+\frac{5}{7}(0) \\ -2=A(-7)+0 \\ -2=A(-7)\]

OpenStudy (anonymous):

so 2/7 ok, ok then, so the value you plug in is always the negative or positive of like (x+6) which goes to (-6+6)

OpenStudy (freckles):

i plugged in -6 becacuse that made x+6 zero just like before when I plugged in 1, that would make x-1 zero

OpenStudy (freckles):

ust like on the previous problem i plugged in 1 because it made two terms zero then i plugged in 0 because it made one term zero

OpenStudy (anonymous):

ok i see that makes the value searching for X or whatever variable you have easier

OpenStudy (anonymous):

so after that you would do u substitution on those answers?

OpenStudy (freckles):

\[\frac{x+4}{(x+6)(x-1)}=\frac{A}{x+6}+\frac{B}{x-1}=\frac{\frac{2}{7}}{x+6}+\frac{\frac{5}{7}}{x-1}\]

OpenStudy (freckles):

then you would integrate that

OpenStudy (anonymous):

1/7lnl(x+6)^2(x-1)^5l +C this is the answer

OpenStudy (freckles):

\[\frac{2}{7} \int\limits_{}^{} \frac{1}{x+6} dx+\frac{5}{7} \int\limits \frac{1}{x-1} dx\]

OpenStudy (freckles):

the factored out 1/7 and use power rule for log and then product rule

OpenStudy (freckles):

you will find the following useful in showing their answer ab+ac=a(b+c) \[r \ln(x)=\ln(x^r) \text{ power rule } \\ \ln(a b)=\ln(a)+\ln(b) \text{ product rule }\]

OpenStudy (anonymous):

so then the 2 in 2/7 would become the power for (x+6) because of the power rule for ln

OpenStudy (freckles):

what do you have so far?

OpenStudy (anonymous):

ln(x+6)^2(x-1)^5

OpenStudy (freckles):

\[\frac{2}{7} \int\limits\limits_{}^{} \frac{1}{x+6} dx+\frac{5}{7} \int\limits\limits \frac{1}{x-1} dx \\ \frac{2}{7} \ln|x+6|+\frac{5}{7}\ln|x-1| +C \\ \frac{1}{7}(2 \ln|x+6|+5 \ln|x-1|)+C \text{ by factoring } \\ \frac{1}{7}(\ln((x+6)^2)+\ln((x-1)^5))+C \text{ by power rule } \\ \frac{1}{7} (\ln((x+6)^2(x-1)^5)+C \text{ by product rule }\]

OpenStudy (anonymous):

thank you again i do have one more question to post